OpenStudy (anonymous):

if 3a-9=6, what is the value of 4a-1? A) -5 B) 19 C) 15 D) 5

OpenStudy (anonymous):

3a - 9 = 6 3a = 9+6 a = (9+6) / 3 put value of a in 4a -1